OpenStudy (anonymous):

Taxi cab charges a $10.75 pickup fee and $1.50 per mile traveled. my total fare for an out of town cab ride is $75.25. how many miles did i travel in the cab?

OpenStudy (ganpat):

analyse the question and get a equation from it.. can u try ??

OpenStudy (ganpat):

10.75 + 1.50x = 72.25.. can u calculate for x ??

OpenStudy (ganpat):

where x is ur miles..

OpenStudy (ganpat):

10.75 + 1.50x = 72.25 1.50x = 72.25 - 10.75 1.50x = 61.5 x = 41 miles...

OpenStudy (anonymous):

ok i see, never been good with word problems. i can calculate the for x.

OpenStudy (ganpat):

ya sure.. u can give a try for next one :)